Directions: Give the total amount of the mix that results and the percent concentration or worth of the mix.
1. Two quarts of 100% orange juice are mixed with 2 quarts of water (0% juice).
So far I have:
(1st quantity*concentration)+(2nd quantity*concentration)=1st+2nd quantity* desired concentration
which I turned into:
2(1)+2(0)=x, where x is the amount of mix...maybe?
I'm not sure if I'm doing this right. My answer is x=2. I'm not even sure if that answers the question! So I am also asking what is the question asking for, besides how do you solve it?

Answer by chessace(471)

50% concentration is the desired answer.
You computed x = # quarts (out of 4 total) of OJ in the final.
Actually, your "x" changed meaning in midstream.
Better continuation of what you wrote:
2*1 + 2*0 = x quantity * y concentration
x = 2+2 (the total volume)
2 = 4 y
y = 1/2 concentration.